# Unraveling the Truth: Do LLMs really Understand Charts? A Deep Dive into Consistency and Robustness

### v1

@@ Abstract
Unraveling the Truth: Do LLMs really Understand Charts?
@@ A Deep Dive into Consistency and Robustness
Srija Mukhopadhyay*, Adnan Qidwai*, Aparna Garimella, Pritika Ramu Vivek Gupta§, Dan Roth§ *IIIT Hyderabad, Adobe Research, §University of Pennsylvania {srija.mukhopadhyay@research, adnan.qidwai@students}.iiit.ac.in, {garimell,pramu}@adobe.com ; {gvivek, danroth}@seas.upenn.edu
footnotetext:  * contributed equally, ‡  corresponding author
@@ 1 Introduction
How do models perform on specific aspects of chart understanding, such as question complexity and chart type? E.g. in Figure 2, the performance of the model is examined across different types of charts—specifically, Simple Charts and Complex Charts—as well as varying question types, including Simple Questions and Complex Questions. Complex Charts, such as grouped-bar charts that compare multiple attributes side by side present information in a more intricate manner compared to Simple Charts, which depict data about a single attribute using a single bar. Similarly, questions can range from complex tasks like identifying maximum values, performing aggregations, and making comparisons, to simpler queries focused on straightforward value extraction. Investigating how models handle these varied chart types and question complexities provides crucial insights into their performance and adaptability.
@@
Furthermore, is the robustness of these models, their ability to generalize across diverse variations, adequately explored? The same dataset can be depicted in multiple visual formats. For instance, Figure 3 demonstrates how an original chart can be transformed into stair plots, bar charts, stacked representations, and many more. These variations can differ in aspects such as color schemes, patterns, legend positioning, and even details specific to each chart type like legend orientation and grid sizes on the x-axis and y-axis. Exploring the effect of these variations could provide deeper insights into the data and enhance the comprehensibility of the visualizations for models.
To answer these questions, we present a rigorous evaluation of leading CQA models on a meticulously curated dataset. This dataset encompasses diverse chart types and question categories, allowing for a thorough assessment of model performance across varying levels of complexity. We examine how well the models generalize across diverse visual representations of identical data, assessing their robustness against perturbations. Our findings reveal significant performance discrepancies, particularly when transitioning from simple to complex chart-question combinations. Moreover, we demonstrate that even the highest-performing models exhibit a substantial drop in accuracy when subjected to diverse perturbations, highlighting the critical need for improved robustness in CQA. This paper makes the following contributions:
@@
Our research sheds light on the current state of CQA, offering crucial insights. Code and data are made publicly available at https://vgupta123.github.io/chartrobustness.html.
@@ 2.1 Dataset Selection
ChartQA includes two distinct question categories: "Human" and "Augmented". "Human" questions were generated by human annotators, while "Augmented" questions were machine-generated, providing a diverse spectrum of question styles. Another important aspect which motivated our choice of ChartQA dataset was the presence of underlying tables. This feature enabled us to generate controlled visual perturbations for the later section of our study. Our experiments were conducted exclusively on the test set of ChartQA, comprising questions, charts and the corresponding tables.
@@ 2.2 Chart and Question Labelling
Charts were classified into two categories using a code-based approach.
- Simple Charts: These charts represent a single entity over a dataframe with two columns, exhibiting no overlaps or complex visual elements. Figure 2 shows an example of such chart titled "Number of Votes given by Area".
- Complex Charts: These charts feature more than two columns, often encompassing multiple entities, leading to increased visual complexity. Figure 2 shows an example of such chart titled "Number of Votes given to Parties by Area".
@@
- Simple Questions: These questions primarily focus on data extraction, and typically involve a single step of reasoning. Figure 2 shows an example of such questions "What is the number of votes given in La La La Land?".
- Complex Questions: These questions require multi-step reasoning and data extraction, often involving comparisons and logical inferences. Figure 2 shows an example of such questions "Is the mean number of ‘Party A’ voters greater than the mean number of ‘Party B’ voters?".
We introduced these categorizations while preserving the existing division of question generation types (human-generated and augmented questions), resulting in eight categories. The number of unique question-chart pairs in each category is presented in Table 1. This detailed categorization allows us to isolate the impact of chart and question complexity on model performance, providing a deeper understanding of their capabilities and limitations.
@@ 3 Experiments
Models
To rigorously assess the performance of CQA models, we selected a diverse range of state-of-the-art models, varying in architecture, size, and training setup. All models were evaluated using a zero-shot Chain-of-Thought Wei et al. (2022) prompting approach, with prompts tailored for each model to maximize performance. Importantly, no additional reasoning aids were provided to any of the models. For the sake of clarity and analysis, we grouped the models into three broad categories:
@@
Mod. Chart-based VLMs Generalist VLMs MLLMs
MatCha UniChart
@@ 1.5 Flash
Table 2: Model performance across different categories. S denotes ’Simple’ and C denotes ’Complex’. The first and second letter represents chart and question type respectively.
Evaluation
To improve on the Relaxed Accuracy metric, we introduce a new evaluation metric that includes extra checks for precise answer matching. This metric, similar to Relaxed Accuracy, provides a 5% leverage for numerical answer matching. However, it includes the following checks:
  • Alphanumeric String Matching: Removing comma and spaces from the during answer matching to ensure an exact alphanumeric string comparison.
  • Unordered Exact List Matching: For questions requiring multiple answers, an unordered exact list matching is applied, which ensure model correctly identifies all the elements in answer list, regardless of their order.
To validate the accuracy of our proposed evaluation metric, we manually verified the answers obtained using this metric.
Smaller VLMs.
Smaller models (QwenVL, CogAgent, InternLM) struggled to produce answers in the correct format. We addressed this by employing an "LLM as an Extractor" approach, using Gemini 1.5 Flash to extract answers from their outputs. Manual verification of 150 samples confirmed that Gemini primarily acted as a formatting tool, preserving the original model’s answer in 149 cases and performing rounding in the one remaining instance. This demonstrates Gemini’s effectiveness in enhancing the usability of smaller models without significantly altering their intent.
4 Can VLMs reasons consistently?
4.1 Results and Discussion
Table 2 gives an overview of all results obtained for this section.
@@ 4.1 Results and Discussion
Interestingly, models specifically trained on chart tasks (MatCha, UniChart) excel on augmented questions. This likely stems from their exposure to similar question formats during training. This is particularly evident in simple questions from the augmented set, where MatCha achieves a high accuracy of 91.40%, followed by UniChart at 87.20%. However, they struggle significantly with reasoning-based questions, achieving as low as 25% accuracy for complex chart and complex question pairs, highlighting the need for enhancement in the reasoning abilities of such models.
@@
Across all models, a consistent trend emerged: performance was consistently better on simple charts compared to complex charts, regardless of the question type. This behavior is likely attributable to the inherent difficulty in understanding and extracting values from complex charts. Factors like overlapping data points and complex color resolution contributes to challenges in data extraction, increasing the difficulty of reasoning on such charts.
@@
For the same chart type, models consistently perform better on simple questions compared to complex questions. This significant difference in scores highlights the limitations of certain models in fine-grained data extraction and reasoning. GPT-4o and Gemini 1.5 Flash exhibit the smallest decrease in scores, indicating strong data extraction and reasoning capabilities. Smaller models, particularly those specifically trained on charts, struggle with questions requiring mathematical reasoning, despite their competence in basic data extraction.
@@
To assess model capabilities, we compared performance on two categories: "Simple Charts, Complex Questions" and "Complex Charts, Simple Questions." This analysis reveals whether a model excels at visual data extraction (complex charts) or reasoning (complex questions).
Our results show that LLMs like GPT-4 demonstrate strong reasoning skills, excelling on complex questions even with simple charts. Conversely, Gemini 1.5 Flash performs consistently across both categories. Generalist and chart-based VLMs tend to favor complex charts over complex questions, suggesting limitations in complex reasoning. This insight allows for targeted model fine-tuning to enhance specific domains where they lack dexterity.
@@
We focused on identifying patterns of model failure across all categories. Given below are a few recurring difficulties for models:
- Charts containing similar colours: Models struggled with charts which required discrimination between slightly different colors. The issue extends further to recognizing specific colors by their names accurately.
- Tight pie charts: In some instances, models incorrectly assigned labels to categories in pie charts with narrow slices. Thus, failing to identify the correct association.
- Charts containing summary statistics: Models failed to interpret such charts, recalculating metrics like mean or sum even though these values were explicitly provided within the chart itself.
- Questions involving counting: Models consistently struggled to accurately count objects when the number exceeded ten.
@@
Surprisingly, even when presented with irrelevant or blank images, some models successfully answered the questions, indicating a reliance on their pre-existing knowledge. This observation suggests potential leaks in testing data, as models even provided factually incorrect answers, highlighting the need for masked evaluation sets for visual reasoning tasks.
@@ 5 Are VLMs robust on CQA?
5.1 Our RobustCQA Dataset
Creation
  • Color Scheme Changes: Modifying color palettes, gradients and hues.
A detailed section showcasing all perturbation types has been presented in the Appendix.
The perturbed charts were generated using MatplotLib, ensuring that only the perturbed element changed while maintaining consistency in all other chart elements. The tables from the ChartQA dataset served as the source for the underlying data.
Human Verification
To ensure the quality and relevance of our dataset, a rigorous manual annotation process was employed. Expert evaluators meticulously verified each perturbed chart, assessing its comprehensibility and answerability by humans. They also evaluated the relevance of each perturbation to the specific chart type, refining the perturbation set to include only meaningful variations. The underlying tables were also thoroughly verified to confirm that the generated questions remained answerable based on the chart data. This comprehensive evaluation was facilitated by a custom-built annotation platform, specifically designed to streamline the manual annotation process and ensure high-quality data.
@@ 1.5 Flash
Table 4: Model Performance on various perturbations on Complex Charts.
Final Dataset.
The finalized perturbations were then grouped into related categories to create the final dataset. This set comprises 22 unique perturbations categories for simple charts and 25 such categories for complex charts, encompassing a wide range of visual variations.
To ensure a fair analysis of model robustness across perturbations, 100 questions were sampled for each question type within each chart category. This standardized question set allows for direct comparison of model performance across different visual representations. A detailed breakdown of the perturbation categories along with examples has been included in the Appendix.
@@ 5.2 Methodology
To delve deeper into the performance and limitations of leading chart question answering models, we evaluated Qwen-VL, CogAgent-VQA, InternLM-XComposer2 (open-source VLMs) and Gemini 1.5 Flash, GPT-4o (closed-source MLLMs) using our RobustCQA dataset. We employed a similar evaluation metric as before, leveraging an LLM extractor for smaller models to ensure consistent output format, and analyzed all models through Zero-Shot Chain-of-Thought prompting.
@@ 5.3 Results and Discussion
The results obtained for perturbations on complex charts have been highlighted in table 4. For simple charts, a similar table has been presented in the Appendix.
@@
The results reveal a significant performance degradation for most models when confronted with perturbations. While performance generally decreases across all models, some exhibit more drastic drops.
Among open-source models, InternLM-XComposer2 proves most resilient, demonstrating consistency across various perturbations. However, CogAgent-VQA and Qwen-VL struggle significantly with most perturbations, exhibiting low accuracy. Surprisingly, GPT-4o, a closed-source model, displays relatively low accuracy with most perturbations, highlighting a potential lack of robust data extraction skills, particularly with non-annotated charts. In contrast, Gemini 1.5 Flash demonstrates notable consistency with minimal variations across perturbations, showcasing strong reasoning and data extraction capabilities.
Manual analysis of model performances suggests that Gemini 1.5 Flash’s success stems from its approximation skills, allowing it to accurately estimate values even in non-annotated charts where other models struggle. This highlights the importance of improving data extraction for non-annotated charts to enhance model robustness in chart-based tasks.
@@
$(Q2)$ Are there certain perturbations which help enhance the model performance?
@@
Our experiments highlighted several perturbations that unexpectedly improved model performance. Notably, across all models, annotated data points consistently boosted accuracy. While the most beneficial plot type varied across models and question/chart categories, annotated bar graphs emerged as a consistently positive influence.
Furthermore, the addition of a grid and altering tick orientation also yielded significant performance improvements. Grids provide models with precise reference points for data estimation, while clear tick labels enhance accurate data point interpretation. Other beneficial perturbations included replacing legends with labels on lines and adjusting legend positioning. Replacing legends reduces the complexity of color resolution, which often leads to accuracy drops.
Optimal legend placement ensures that labels do not obscure crucial data points. Initial experimentation also indicated that increasing font size, particularly for smaller models, often led to improved performance, results for which have been presented in the Table 5.
$(Q3)$ Are there perturbations which are always detrimental to model performance?
@@
Our analysis reveals that while models demonstrate promising performance on standard chart datasets, they struggle with robustness when faced with visual perturbations. While annotations generally improve performance, most other perturbations negatively impact model accuracy.
Logarithmic scales pose the most significant challenge, likely due to the inherent difficulty in data retrieval even for humans. Models also struggle significantly with horizontal chart variations, particularly horizontal stacked charts, though they perform better with vertical stacked charts.
It should however be noted that models do also struggle with normal stacked plots as well as area plots. This observation can be attributed to the complex process of accurate data extraction for these plots, requiring additional mathematical reasoning to get the precise value of a point which is challenging for humans as well.
Stair plots, where models struggle to identify the precise data point to refer to, and changes in chart scale, which seem to disrupt model attention, also contribute to performance decline.
These findings emphasize the need to develop more robust models that can effectively interpret visual information beyond simple visual cues. Some examples are mentioned in 6.

### v2

# Unraveling the Truth: Do VLMs really Understand Charts? A Deep Dive into Consistency and Robustness

@@ Abstract
Unraveling the Truth: Do VLMs really Understand Charts?
@@ A Deep Dive into Consistency and Robustness
Srija Mukhopadhyay*, Adnan Qidwai*, Aparna Garimella, Pritika Ramu Vivek Gupta, Dan Roth§ *IIIT Hyderabad, Adobe Research, Arizona State University, §University of Pennsylvania {srija.mukhopadhyay@research, adnan.qidwai@students}.iiit.ac.in, {garimell,pramu}@adobe.com; vgupt140@asu.edu; danroth@seas.upenn.edu
footnotetext:  * contributed equally, ‡ corresponding author (work done at UPenn)
@@ 1 Introduction
How do models perform on specific aspects of chart understanding, such as question complexity and chart type? Figure 2 shows the different types of charts across which the performance of a model can be evaluated—specifically, Simple Charts and Complex Charts—along with the different possible question types, including Simple Questions and Complex Questions. Complex Charts, such as grouped-bar charts that compare multiple attributes side by side present information in a more intricate manner compared to Simple Charts, which depict data about a single attribute using a single bar. Similarly, questions can range from complex tasks like identifying maximum values, performing aggregations, and making comparisons, to simpler queries focused on straightforward value extraction. Investigating how models handle these varied chart types and question complexities provides crucial insights into their performance and adaptability.
@@
Furthermore, is the robustness of these models, their ability to generalize across diverse variations, adequately explored? The same data can be depicted in multiple visual formats. For instance, Figure 3 demonstrates how an original chart can be transformed into stair plots, bar charts, stacked representations, and many more. These variations can differ in aspects such as color schemes, patterns, legend positioning, and even details specific to each chart type like legend orientation and grid sizes on the x-axis and y-axis. Exploring the effect of these variations could provide deeper insights into the data and enhance the comprehensibility of the visualizations for models.
To answer these questions, we present a rigorous evaluation of leading CQA models on a meticulously curated dataset. This dataset includes diverse chart types and question categories, allowing for a thorough assessment of model performance across varying levels of complexity. We examine how well the models generalize across diverse visual representations of identical data, assessing their robustness against perturbations.
Our findings reveal significant performance discrepancies, particularly when transitioning from simple to complex chart-question combinations. Moreover, we demonstrate that even the highest-performing models exhibit a substantial drop in accuracy when subjected to diverse perturbations, highlighting the critical need for improved robustness in CQA. This paper makes the following contributions:
@@
Our research sheds light on the current state of CQA, offering crucial insights. Our datasets, along with all the associated scripts, are available at https://robustcqa.github.io/
@@ 2.1 Dataset Selection
ChartQA includes two distinct question categories: "Human" and "Augmented". "Human" questions were generated by human annotators, while "Augmented" questions were machine-generated, ensuring a diverse spectrum of question styles. Another important aspect which motivated our choice of ChartQA dataset was the presence of underlying tables. This feature enabled us to generate controlled visual perturbations for the later section of our study. Our experiments were conducted exclusively on the test set of ChartQA, comprising questions, charts and the corresponding tables.
@@ 2.2 Chart and Question Labelling
The tables provided by ChartQA were loaded as a pandas dataframe. We classify charts as either simple or complex, based on column count in the dataframe: two columns indicate a simple chart, while more than two columns signify a more complex chart. We leverage this fact to classify them using a python script.
- Simple Charts: The tables of these charts contain two columns to represent the dependent and independent variable. The charts thus formed represent a single entity and exhibit no overlaps or complex visual elements. Figure 2 shows an example of such chart titled "Number of Votes given by Area".
- Complex Charts: The tables of these charts feature more than two columns, often encompassing multiple dependent variables. Thus, the charts formed have increased visual complexity. These charts usually depict multiple entities over a common series. Figure 2 shows an example of such chart titled "Number of Votes given to Parties by Area".
@@
- Simple Questions: These questions primarily focus on data extraction, and typically involve a single step of reasoning. A human annotator can ideally answer such a question in a single step. Figure 2 shows an example of such questions "What is the number of votes given in La La La Land?". One can simply answer the question by fetching the value from the chart.
- Complex Questions: These questions require multi-step reasoning along with data extraction, and often involve comparisons and logical inferences. If it takes multiple steps for the human annotator to answer a question, it would be classified as a complex question. Figure 2 shows an example of such questions "Is the mean number of ‘Party A’ voters greater than the mean number of ‘Party B’ voters?". For this question, one would require multiple calculations to reach the final answer.
We introduced these categorizations while preserving the existing division of question generation types (human-generated and augmented questions) which was present in the original dataset, resulting in eight categories. The number of unique question-chart pairs in each category is presented in Table 1. We call this modified dataset ChartQA-Split. The detailed catagorization in our dataset allows us to isolate the impact of chart and question complexity on model performance, providing a deeper understanding of their capabilities and limitations.
@@ 3 Experiments
Models.
To rigorously assess the performance of CQA models, we selected a diverse range of state-of-the-art models, varying in architecture, size, and training setup. All models were evaluated using a zero-shot Chain-of-Thought Wei et al. (2022) prompting approach. An example of our prompt can be found in Figure 4. It is important to note that no additional reasoning aids were provided to any of the models. For the sake of clarity and analysis, we grouped the models into three broad categories:
@@
Evaluation
To evaluate our models, we decided to utilize the Relaxed Accuracy metric owing to the objective nature of the expected answers. To improve on the Relaxed Accuracy metric, we introduce extra checks for precise and accurate answer matching. This metric, similar to Relaxed Accuracy, provides a 5% leverage for numerical answer matching. However, it includes the following checks:
Alphanumeric String Matching: Removing comma and spaces from the given answer and gold label to ensure an exact alphanumeric string comparison.
Unordered Exact List Matching: For questions requiring multiple answers, an unordered exact list matching is applied, to ensure that the model correctly identifies all the expected elements in answer list, regardless of their order.
Furthermore, to validate the accuracy of our proposed evaluation metric, we manually verified the answers obtained using this metric. Our metric is usable and applicable for general large-scale model evaluation in question-answering based tasks.
Smaller VLMs.
We noticed that smaller models (QwenVL, CogAgent, InternLM) struggled to produce answers in the correct format. This might be due lack of complex instruction following abilities. We addressed this by using Gemini 1.5 Flash to extract answers from their outputs in a favourable format, hence using the LLM as an extractor. Manual verification of 150 samples confirmed that Gemini 1.5 Flash primarily acted as a formatting tool, preserving the original model’s answer in 149 cases and performing rounding in the one remaining instance. This demonstrates Gemini’s effectiveness in enhancing the usability of smaller models without significantly altering their intent. The prompt used, can be found in Figure 5.
4 Can models reasons consistently?
4.1 Results and Discussion
Table 2 gives an overview of all results obtained for this section.
Type Chart-based VLMs Generalist VLMs MLLMs
Mat-
Cha
Uni-
Chart
@@ 1.5 Flash
Table 2: Model accuracy across different categories. S denotes ’Simple’ and C denotes ’Complex’. The first and second letter represents chart and question type respectively.
@@
Models specifically trained on chart reasoning tasks (MatCha, UniChart) show exceptional performance while answering augmented questions, as highlighted in previous work as well (Liu et al. (2023b); Masry et al. (2023). This likely stems from their exposure to similar question formats during training, which is particularly evident in simple questions from the augmented set, where MatCha achieves a high accuracy of 91.40%, followed by UniChart at 87.20%. However, they struggle significantly with reasoning-based questions, achieving as low as 25% accuracy for complex chart and complex question pairs, highlighting the need for enhancement in the reasoning abilities of such models.
@@
Across all models, a consistent trend emerged: performance was consistently better on simple charts compared to complex charts, while comparing with the same question type. This behavior is likely attributable to the inherent difficulty in understanding and extracting values from complex charts. Factors like overlapping data points and the requirement of precise color resolution contributes to challenges in data extraction, increasing the difficulty of reasoning on such charts.
@@
For the same chart type, models consistently perform better on simple questions compared to complex questions. This significant difference in scores highlights the limitations of certain models in fine-grained data extraction and reasoning. GPT-4o and Gemini 1.5 Flash exhibit the smallest decrease in scores, indicating strong reasoning capabilities along with commendable data extraction skills. Smaller models, particularly those specifically trained on charts, struggle with questions requiring mathematical reasoning, despite their competence in basic data extraction.
@@
To further assess model capabilities, we compared performances of models on two categories: "Simple Charts, Complex Questions" and "Complex Charts, Simple Questions." This analysis reveals whether a model excels at visual data extraction (complex charts) or reasoning (complex questions).
Our results show that MLLMs like GPT-4o demonstrate strong reasoning skills, excelling in answering complex questions. Gemini 1.5 Flash on the other hand performs consistently across both categories. Generalist and chart-based VLMs tend to favor the complex chart, simple question pair over the simple chart, complex question pair, suggesting limitations in reasoning abilities. This insight allows for targeted model fine-tuning to enhance specific domains where they lack dexterity.
@@
We focused on identifying patterns of model failure across all categories. In total, we found that 181 questions could not be answered by any model that we tested on. Given below are a few recurring difficulties that models faced:
36/181 - Questions involving counting: Models consistently struggled to accurately count objects when the number exceeded ten.
30/181 - Charts containing similar colours: Models struggled with charts which required discrimination between highly similar colors or shades of the same color.
17/181 - Identifying colours from name: Models struggle to accurately identify the color of chart elements when prompted to do so.
17/181 - Charts involving summary statistics: Models struggle to interpret charts with summary statistics, often confusing presented values with the need for recalculation. For example, given a chart of "Average of company revenues," they struggle to answer questions about "company A’s average revenue," unsure whether to extract the value or recalculate it. This highlights a key limitation in their understanding of statistical representations.
7/181- Tight pie charts: In some instances, models incorrectly assigned labels to categories in pie charts with narrow slices, hence failing to identify the correct association.
A more detailed analysis on this topic has been presented in the Appendix.
@@
Surprisingly, even when models were presented with irrelevant or blank images, some models successfully answered the questions, indicating a reliance on their pre-existing knowledge. This observation suggests potential leaks in testing data. Models were able to provide the correct answer even if the answer was factually incorrect from a real-world point-of-view, highlighting the need for masked evaluation sets for visual reasoning tasks.
@@
Table 4: Model Performance on various perturbations on Complex Charts
5.1 Our RobustCQA Dataset
Creation
  • Color Scheme Changes: Modifying color palettes, gradients and hues.
The perturbed charts were generated using the Matplotlib library. We ensured that only one element is altered per perturbation keeping the rest of the elements the same. The tables from the ChartQA dataset served as the source for the underlying data.
Human Verification
To ensure the quality and relevance of our dataset, a rigorous manual annotation process was employed. Expert evaluators meticulously verified each perturbed chart, assessing how easily comprehensible and answerable each perturbed chart was. They also evaluated the relevance of each perturbation to the specific chart type, refining the perturbation set to include only meaningful variations. The underlying tables were also thoroughly verified to confirm that the generated questions remained answerable based on the chart data. This comprehensive evaluation was facilitated by a custom-built annotation platform, specifically designed to streamline the manual annotation process and ensure high-quality data.
Final Dataset
The original 75 perturbations were then grouped into categories of related perturbations to create the final dataset. This set consists of 22 unique perturbation categories for simple charts and 25 such categories for complex charts, covering a wide range of visual variations.
To ensure a fair analysis of model robustness across perturbations, 100 questions were sampled for each chart type and question type pair. This resulted in a total of 400 unique table and QA pairs for our final evaluation. This standardized question set allows for direct comparison of model performance across different visual representations. We finally compare the results of all perturbations against the basic or default Matplotlib chart.
A detailed breakdown of the perturbation categories along with examples has been included in the Appendix.
@@ 5.2 Methodology
To delve deeper into the performance and limitations of leading chart question answering models, we evaluated Qwen-VL, CogAgent-VQA, InternLM-XComposer2 (open-source VLMs) and Gemini 1.5 Flash, GPT-4o (closed-source MLLMs) using our RobustCQA dataset. We employed a similar evaluation metric as described previously, leveraging an extractor LLM for smaller models to ensure consistent output format, and analyzed all models through Zero-Shot Chain-of-Thought prompting.
@@ 5.3 Results and Discussion
The results obtained for perturbations on complex charts have been highlighted in table 4. The results for perturbations on simple charts has been presented in table 9 in the Appendix.
@@
The results reveal a significant performance degradation for most models when confronted with perturbations. While performance generally decreases across all models, some exhibit more drastic drops. Among open sourced models, InternLM-XComposer2, and among closed source models, Gemini 1.5 Flash proved to be the most consistent across various perturbations. Open sourced models like CogAgent-VQA and Qwen-VL and even closed source models like GPT-4o displayed relatively low accuracy with most perturbations, highlighting a potential lack of robust data extraction skills. Manual analysis of the responses from the models highlighted the importance of improving data extraction for non-annotated charts to enhance model robustness in chart-based tasks.
@@
$(Q2)$ Are there specific perturbations that help enhance model performance?
@@
Our experiments highlighted several perturbations that improved model performance. Across all models, annotated data points consistently boosted accuracy. While the most beneficial plot type varied across models and question/chart categories, annotated bar graphs emerged as a consistently positive influence.
In addition to that, grids to act as reference points for data extraction and better tick orientation also contributed positively. Furthermore, labelling the lines to reduce the complexity of color resolution and placing the legend optimally to ensure that it doesn’t obscure crucial data points also helped the models. We also noticed that increasing the font size played an important role in aiding all models, especially smaller models. The results for the same have been presented in Table 5.
$(Q3)$ Are there perturbations which are always detrimental to the model performance?
@@
Our analysis reveals that while models demonstrate promising performance on standard chart datasets, they struggle with robustness when faced with visual perturbations. While annotations generally help improve model performance, most other perturbations negatively impacted model accuracy.
Notable ones among them include logarithmic scales which can be challenging for humans as well. Additionally, models also struggle significantly with horizontal chart variations, particularly horizontal stacked charts. In general it was noticed that models struggled to reason on stacked charts, possibly due to the requirement of additional mathematical reasoning for data extraction. Stair plots also caused significant trouble as models could not identify the precise data point to refer to. Our findings emphasize the need of more diverse datasets along with more robust models that can effectively interpret visual information beyond just simple visual cues.
Our analysis suggests chart type effectiveness varies by question type. For instance, line charts help in visualizing trends and correlations. Stacked bar charts are generally unsuitable except for questions that require data aggregation. Bar charts, while useful for comparing individual values within a certain group, prove to not be good for showcasing correlations across different groups or entities. Our analysis helps with understanding and creating suitable charts for domain specific tasks.
$(Q5)$ Does the effect of each perturbation type vary across models?
The impact of each perturbation on model performance exhibits significant variation. While question and chart type play a role, for a given model, certain perturbations consistently prove more helpful or harmful. This nuanced effect of perturbation type on the model performance is detailed in Table 7, 8. We believe that our analysis allows us to identify specific areas for helping improve each model through targeted model fine-tuning.
Additional insights and details obtained from our experiments have been presented in the Appendix.
6 Related Work
Chart comprehension and question answering (CQA) are critical domains with a growing body of research. While existing CQA datasets assess models’ advanced reasoning capabilities over charts, many face significant limitations. These include small dataset sizes Kim et al. (2020), reliance on template-based questions and synthetically generated charts Methani et al. (2020); Chaudhry et al. (2020); Kafle et al. (2018); Han et al. (2023), restriction to specific domains Methani et al. (2020); Ahmed et al. (2023); Li and Tajbakhsh (2023), or focusing solely on open-domain question answering Kantharaj et al. (2022). Even the current state-of-the-art dataset, Chart QA Masry et al. (2022), has limitations due to the lack of classification labels for more meaningful analysis, and limited variation in chart types.
More recent datasets, such as ChartX Xia et al. (2024), have expanded the range of chart types analyzed. ChartBench Xu et al. (2024) and MMC Liu et al. (2024) focus on large-scale datasets with more diverse chart types.
A very recent work, CharXiv Wang et al. (2024), provides extensive evaluations across a range of charts and questions, including both reasoning-based and descriptive queries. They also perform ablation studies by modifying charts and questions.
However, to the best of our knowledge, RobustCQA is the first dataset to systematically perturb all elements within a chart, enabling fine-grained analysis of factors affecting model performance. Additionally, we conduct a detailed analysis of the Chart QA dataset based on question and chart complexity, which has not been done at this level of detail before.
Modeling approaches for charts
Various approaches have been developed for chart modeling. This includes models specifically designed for chart comprehension and reasoning, built with the end-to-end goal of reasoning over charts Liu et al. (2023b); Masry et al. (2023); Singh and Shekhar (2020), as well as models that convert charts into intermediate table formats Liu et al. (2023a), enabling reasoning by generalized large language models (LLMs) through Chain of Thought Wei et al. (2022) or Program of Thought Chen et al. (2023) prompting. Additionally, there are generalized models used for multi-modal reasoning tasks, including chart comprehension Team et al. (2023); Achiam et al. (2023); Bai et al. (2023b); Dong et al. (2024); Hong et al. (2024). Recent efforts have also focused on developing smaller, yet accurate models for this task Wang et al. (2024).
While these approaches have shown significant progress, their specific failure points remain unclear. A recent study Islam et al. (2024) analyzed the performance of GPT-4v and Gemini, providing a broad evaluation of these models across various chart comprehension tasks, including question answering, summarization, and fact-checking. In contrast, our work focuses specifically on CQA across a broader range of models, offering an in-depth analysis of the question and chart types contributing to model failures. Our contribution identifies the exact question types and chart elements that lead to model errors, offering key insights to improve model performance.
Vision-Language Model Robustness
Recent studies have highlighted the vulnerability of models to attacks and perturbations Ma et al. (2024); Zhao et al. (2023), raising concerns about their robustness in real-world deployment. Motivated by this, we developed a robustness benchmark specifically for chart question answering (CQA). While previous work Gupta et al. (2024) analyzed models like DePlot and MatCha on perturbed charts, focusing on questions related to structural and visual context, our study examines general reasoning questions. This approach helps us assess how variations in the visual representation of the same data affect model performance.
7 Conclusion
Future Directions.
Our perturbation analysis provides a nuanced understanding of model performance by revealing both universal and model-specific vulnerabilities and strengths. This insight drives targeted improvements: Model Pretraining: Focusing on perturbations that affect models allows for effective fine-tuning to address weaknesses. Perturbation-Aware Training: Integrating specific perturbations during training enhances overall robustness, helping models develop resilience against challenges. Interpretable Models: Understanding the impact of perturbations aids in debugging and building explainable models, fostering the development of reliable and transparent chart understanding and reasoning systems.
Limitations
The presented work exhibits a few limitations. First, our data was obtained from a singular dataset, and we used only one plotting software for testing the perturbations. Expanding the dataset to include diverse sources and exploring various plotting libraries would strengthen the findings and improve generalizability. Second, the dataset is limited to English, while models are developed and evaluated on a wide variety of languages. Future research is required to expand the domain beyond English. Third, we were not able to cover a few chart types in the course of our analysis in order to make a more generalized perturbation set. This included pie and doughnut charts, pyramid and funnel charts as well as radar charts. Due to metadata limitations and the complexity of adapting data for chart representation, these charts were excluded. Fourth, inconsistent metadata of the original dataset sometimes lacked visual captions present in the original charts, which could not be replicated in the perturbed charts. Because of this, we were not able to identify attributes pertaining to chart elements, for example, colour.
Ethics Statement
Acknowledgement
Appendix
Effect of Font-size on Models
Table 5 illustrates the significant impact of font size on model performance. Increasing the font size improves the OCR capabilities of visual language models (VLMs), suggesting that larger font sizes can enhance model accuracy in chart comprehension tasks. This finding indicates that adjusting font size could serve as an effective preprocessing step for boosting performance in such tasks.
Gemini 1.5 Flash Qwen VL
Perturbation types Small Font Big Font Small Font Big Font
Normal line plot 62 63 6 27
Colors in a given
scheme (line)
56 63 7 26
Colors random (scatter) 46 57 5 21
Line Represntation 47 50 10 18
Stem Plot 45 52 6 15
Stair Plot 42 48 10 26
Ablation - removing
Y axis
63 64 7 22
Rotated X axis Tick 56 62 8 22
Annotated Bar Graph 77 82 12 42
Horizontal Bar Graph 54 65 11 19
Table 5: Effect of increasing font size
Model Scores
Alongside the scores of models across various perturbations for complex charts provided in Table 4, we have also presented the model scores for different perturbations in simple charts in Table 9.
Where can models not answer?
Continuing with the analysis of Q5 in section 4.1, we provide a further breakdown of cases where models failed to answer correctly, as shown in Table 6. In addition to previously discussed issues, we found that in some instances (4/181), models did not fully comprehend the entire chart before answering.
Type of chart Wrong Total Error Rate
Bar 121 1842 6.56%
Line 44 380 11.57%
Pie 16 151 10.59%
Total 181 2373 7.62%
Table 6: Error distribution among different chart types
For example, in a chart with four columns labeled “No (low confidence), No (high confidence), Yes (low confidence), and Yes (high confidence),” models were asked to calculate the percentage of ’No’. However, they failed to recognize that the question required summing the percentages from both ‘No’ columns and instead provided the percentage from only one column. This error highlights a potential gap in the models’ ability to fully integrate image encoding with language decoding, suggesting improvements could be made to better interpret such visual data.